# State space representations of the Roesser type for convolutional layers

### v1

@@ 6 Conclusion
In this work, we constructed state space representations for a general class of convolutional layers. For 2-D convolutional layers with $c_{\mathrm{in}}=c_{\mathrm{out}}$, we further showed minimality of our state space realization. In the field of neural network analysis and synthesis based on LMIs, such easy-to-compute and compact state representations make convolutional neural networks amenable to LMI-based analysis with improved scalability. To this end, state space representations are advantageous over other reformulations of convolutional layers such as reformulations as fully-connected layers due to their compactness. The next step is to find minimal state space representations for all input and output channel sizes. Additional future work includes the Lipschitz constant estimation for general convolutional neural networks and the synthesis of Lipschitz-bounded convolutional neural networks using the state space representation established in this work.

### v2

@@ 6 Conclusion
In this work, we constructed state space representations for a general class of convolutional layers. For 2-D convolutional layers with $c_{\mathrm{in}}=c_{\mathrm{out}}$, we further showed minimality of our state space realization. In the field of neural network analysis and synthesis based on LMIs, such easy-to-compute and compact state representations make convolutional neural networks amenable to LMI-based analysis with improved scalability. To this end, state space representations are advantageous over other reformulations of convolutional layers such as reformulations as fully-connected layers due to their compactness. The next step is to find minimal state space representations for all input and output channel sizes. Additional future work includes Lipschitz constant estimation for general convolutional neural networks and the synthesis of Lipschitz-bounded convolutional neural networks using the state space representation established in this work.